find the equation of slope=2 and x-intercept=9

Answer by nerdybill(7384)

find the equation of slope=2 and x-intercept=9
.
Essentially, the point where you have the x-intercept=9 is at (9,0)
Plug the point and the slope into the "point-slope" form:
y - y1 = m(x - x1)
y - 0 = 2(x - 9)
y = 2(x - 9)
y = 2x - 18 (this is what they're looking for in "slope-intercept" form)